A valószínűségszámítás gyakorlaton való részvétel

dátum:  II. 7. II. 14. II. 21. II. 28. III. 7. III. 14. III. 21.

diákok:

Angyal Dóra  +  −  +  +  +  +  +
Arnót Titanilla  −  +  +  −  −  −  −
Bakó Ágnes  +  +  +  +  +  +  +
Bathó Gábor  +  −  +  −  +  +  +
Bíró Beáta  −  −  −  −  −  −  −
Bócsi Judit  +  −  +  +  +  +  −
Boda Judit  +  +  +  +  +  +  +
Egyed Károly  −  +  −  −  +  +  +
Ficzere Éva  −  +  −  −  +  +  +
Fista Norbert  +  +  +  +  +  +  −
Gyöngyösi Levente  +  +  +  −  +  +  −
Helmeczi József  −  +  +  +  −  −  −
Jónás Ágnes  +  +  +  +  +  +  −
Karácsony Szilveszter  +  +  +  +  −  +  +
Kalóz Margit  +  +  +  −  +  +  +
Kis Mirella  +  +  −  +  +  +  +
Kovács Viktor  +  +  +  +  −  +  +
Ködöböcz József  −  +  −  +  +  +  −
Lengyel Sándor  −  −  +  +  −  +  +
Mályi Nikoletta  +  +  +  +  +  +  +
Miskolczi Panna  −  +  −  +  +  +  +
Nagy Beáta  +  +  +  −  +  +  +
Nagy Eszter  −  +  +  +  +  +  +
Papp Ágnes  +  +  +  +  +  −  +
Pokol László  +  −  −  −  −  +  −
Pósa Csenge  −  +  +  +  +  +  +
Szántó Ágnes  −  +  −  −  −  −  −
Szeifert Diana  −  +  +  +  +  −  +
Szoó Tibor  +  −  −  −  −  −  −
Szpodnyi Marianna  −  −  +  −  −  −  −
Tóth Edit  +  +  +  +  +  +  +
Tóth Veronika  −  −  +  +  −  −  +

dátum:  III. 28. IV. 4. IV. 18. IV. 25. V. 2. V. 9. V. 16.

diákok:

Angyal Dóra  +  −  +  +  +  +
Arnót Titanilla  −  −  −  −  −  −
Bakó Ágnes  +  +  +  +  +  +
Bathó Gábor  +  −  −  +  +  +
Bíró Beáta  −  −  −  −  −  −
Bócsi Judit  +  −  +  −  +  +
Boda Judit  +  +  +  +  +  +
Egyed Károly  +  +  −  +  −  −
Ficzere Éva  +  −  +  +  +  +
Fista Norbert  +  +  +  +  −  −
Gyöngyösi Levente  +  +  +  +  +  +
Helmeczi József  +  +  +  +  −  +
Jónás Ágnes  +  +  −  +  +  +
Kalóz Margit  +  −  +  +  +  +
Karácsony Szilveszter  +  +  +  −  +  +
Kis Mirella  +  +  +  −  +  +
Kovács Viktor  +  +  +  −  −  +
Ködöböcz József  +  −  +  −  +  −
Lengyel Sándor  +  +  −  +  +  +
Mályi Nikoletta  +  +  −  +  +  +
Miskolczi Panna  −  +  +  +  +  +
Nagy Beáta  +  +  −  +  +  +
Nagy Eszter  +  +  +  +  +  +
Papp Ágnes  +  +  +  +  +  +
Pokol László  −  −  −  −  −  −
Pósa Csenge  +  +  +  −  −  −
Szántó Ágnes  −  −  −  −  −  −
Szeifert Diana  +  +  +  +  +  +
Szoó Tibor  −  −  −  −  −  −
Szpodnyi Marianna  −  −  −  −  −  −
Tóth Edit  +  +  +  +  +  +
Tóth Veronika  +  +  +  −  +  +